On January 1, 2018, the Mason Manufacturing Company began construction of a building to be used as its office headquarters. The building was completed on September 30, 2019.

Expenditures on the project were as follows:

January 1, 2018 $ 2,020,000
March 1, 2018 1,740,000
June 30, 2018 1,940,000
October 1, 2018 1,740,000
January 31, 2019 441,000
April 30, 2019 774,000
August 31, 2019 1,071,000


On January 1, 2018, the company obtained a $4,900,000 construction loan with a 12% interest rate. The loan was outstanding all of 2018 and 2019. The company’s other interest-bearing debt included two long-term notes of $2,000,000 and $8,000,000 with interest rates of 8% and 10%, respectively. Both notes were outstanding during all of 2018 and 2019. Interest is paid annually on all debt. The company’s fiscal year-end is December 31.

Required:
1. Calculate the amount of interest that Mason should capitalize in 2018 and 2019 using the specific interest method.
2. What is the total cost of the building?
3. Calculate the amount of interest expense that will appear in the 2018 and 2019 income statements

On January 1, 2018, the Mason Manufacturing Company began construction of a building to be used as its office headquarters. The building was completed on September 30, 2019.

Expenditures on the project were as follows:

January 1, 2018 $ 2,020,000
March 1, 2018 1,740,000
June 30, 2018 1,940,000
October 1, 2018 1,740,000
January 31, 2019 441,000
April 30, 2019 774,000
August 31, 2019 1,071,000


On January 1, 2018, the company obtained a $4,900,000 construction loan with a 12% interest rate. The loan was outstanding all of 2018 and 2019. The company’s other interest-bearing debt included two long-term notes of $2,000,000 and $8,000,000 with interest rates of 8% and 10%, respectively. Both notes were outstanding during all of 2018 and 2019. Interest is paid annually on all debt. The company’s fiscal year-end is December 31.

Required:
1. Calculate the amount of interest that Mason should capitalize in 2018 and 2019 using the specific interest method.
2. What is the total cost of the building?
3. Calculate the amount of interest expense that will appear in the 2018 and 2019 income statements
Answer #1
2018 2019
Interest Capitalized $               5,85,000 $         7,33,752
Interest Expense $               9,63,000 $         8,14,248
Cost of Building = $   1,10,44,752
Workings:
Expenditure for 2018
Jan 1,2018 $            20,20,000 X 12/12 = $        20,20,000
March 1, 2018 $            17,40,000 X 10/12 = $        14,50,000
June 30, 2018 $            19,40,000 X 6/12 = $          9,70,000
October 1, 2018 $            17,40,000 X 3/12 = $          4,35,000
$            74,40,000 $        48,75,000
Interest Capitalized in 2018
$            48,75,000 X 12.00% = $          5,85,000
Expenditure for 2019
Jan 1, 2019 ($74,40,000+$5,85,000) $            80,25,000 X 9/9 = $        80,25,000
Jan 31, 2019 $               4,41,000 X 8/9 = $          3,92,000
April 30, 2019 $               7,74,000 X 5/9 = $          4,30,000
August 31. 2019 $            10,71,000 X 1/9 = $          1,19,000
$         1,03,11,000 $        89,66,000
Interest Capitalized in 2019
$                       89,66,000
Less: $                       49,00,000 X 12.00% X 9/12 = $          4,41,000
$                       40,66,000 X 9.60% X 9/12 = $          2,92,752
Interest Capitalized in 2019 = $          7,33,752
Weighted Average rate of all debt:-
$            20,00,000 X 8% = $          1,60,000
$            80,00,000 X 10% = $          8,00,000
$         1,00,00,000 $          9,60,000
Weighted Average rate of all debt = 9.60%
($9,60,000 / $1,00,00,000)
Interest Expense for 2018:
Total Interest Incurred ($4900000*12%)+($2000000*8%)+($8000000*10%) = $        15,48,000
Less : Interest Capitalized = $          5,85,000
2018 Expense = $          9,63,000
Interest Expense for 2019:
Total Interest Incurred = $        15,48,000
Less : Interest Capitalized = $          7,33,752
2019 Expense = $          8,14,248
Cost of Building
Expenditure for 2018 = $        74,40,000
Interest Capitalized in 2018 = $          5,85,000
Expenditure for 2019 = $        22,86,000
Interest Capitalized in 2019 = $          7,33,752
Cost of Building = $    1,10,44,752